EPJ manuscript No. (will be inserted by the editor)

Insert your title here

Do you have a subtitle? If so, write it here

First author^{1} and Second author^{2}^{a}

1 Insert the first address here 2 the second here

Received: date / Revised version: date Abstract. Insert your abstract here. PACS. PACS-key discribing text of that key – PACS-key discribing text of that key

1 Introduction

Your text comes here. Separate text sections with

2 Section title

and [1]

2.1 Subsection title as required. Don’t forget to give each section and subsection a unique label (see Sect. 2).

Send offprint requests to: a Present address: Insert the address here if needed

Fig. 1. Please write your figure caption here

References

1. Author, Journal Volume, (year) page numbers.

Fig. 2. Please write your figure caption here

Table 1. Please write your table caption here

first  second third
number number number
number number number

2. Author, Book title (Publisher, place year) page numbers